Question Assume the exchange rate for the Mexican Peso is falling relative to the U.S. dollar. An American 17 company will incur losses from this falling exchange rate if the company is making: Not yet answered Select one: Marked out of 1。a. Credit purchases from Mexican companies at prices stated in Mexican Pesos. Flag question b. Credit purchases from Mexican companies at prices stated in U.S, dollars. c.Credit sales to Mexican companies at prices stated in U.S. dollars. O d. Credit sales to Mexican companies at prices stated in Mexican Pesos.
Question 18 Not yet answered a. Affect the income of a prior period. Maked out of 1 Ob.Be associated with a segment of the business that has been discontinued during the current To qualify as an extraordinary item, a gain or loss must: Select one: O b. Be associated with a segment of the business that has been discontinued during the current Flag question eriod O c. Be material in amount, unusual in nature, and not expected to recur. d. Be larger in amount than any other item in the income statement
Question 19 A company had 125,000 shares of common stock outstanding on January 1 and then sold 35,000 additional shares on March 30. Net income for the year was $594,750. What are earnings per share? Select one: O a. $6.61 Not yet answered Marked out of 1 Flag question b. $4.58 ° C. $4.73 O d. $3.93
Question 20 Hines Cannery issued capital stock in 2018 for $700,000. During 2018 the company paid dividends of $250,000. What is the effect of these events in Hines statement of cash flows for 2018? Not yet answered Select one: Marked out o 1 a. $700,000 cash provided by financing activities, and $250,000 cash disbursed for investing activities Flag queston Ob. $700 O b. $700,000 cash provided by financing activities, and $250,000 cash disbursed for operating activities. Oc.$700,000 cash provided by investing activities, and $250,000 cash disbursed for financing activities. O d. $450,000 net cash provided by financing activities.
Answer #1

Answers :

Question: 17) D) Credit Sales to Mexican Companies at prices stated in Mexican Pesos

For US Company Any Transaction Stated in Home Currency (US Dollar) will not result in loss Due to the Exchange rate Fluctuation. But Stated in other Currency ( Peso in question) will result in a gain if purchases made and loss with sales stated in other currency. US Firm will incur a loss when converting the Peso into Dollars.

Question: 18 C) Be Material in amount, Unusual in nature, and not Expected to recur

Extraordinary Gain or loss will Get out of the nonregular course business Activities. Unusual and unexpected in Nature

Question: 19 D) $ 3.93

Earnings Per Share = Net Income / Weighted Average share

Weighted Average Shares

125,000 Share for 12 Months =125000*12 = 1,500,000

35,000 shares for 9 Months = 35,000 *9 = 315,000

weighted average Number of shares = ( 1500,000+315,000)/12 = 151,250 Shares

Earnings Per Share : 594,750 / 151,250 = $3.93223 (or) $3.93 (Answer)

Question :20 D) $ 450,000 net cash Provided by Financing Activities

Issues of share Capital & Dividend Paid Both Activities are financial Activities under Cash Flow Preparation :

Issue of $ 700,000 Capital is the Inflows , Dividend Paid of $250,000 is the Outflows

Net Flows From the Financial Activity Is( $700,000 -$250,000) $ 450,000